# How to Re-enable PDE Loss for Physical Systems Modeling Under Partial Observation

### v1

@@ Abstract
In science and engineering, machine learning techniques are increasingly successful in physical systems modeling (predicting future states of physical systems). Effectively integrating PDE loss as a constraint of system transition can improve the model’s prediction by overcoming generalization issues due to data scarcity, especially when data acquisition is costly. However, in many real-world scenarios, due to sensor limitations, the data we can obtain is often only partial observation, making the calculation of PDE loss seem to be infeasible, as the PDE loss heavily relies on high-resolution states. We carefully study this problem and propose a novel framework named Re-enable PDE Loss under Partial Observation (RPLPO). The key idea is that although enabling PDE loss to constrain system transition solely is infeasible, we can re-enable PDE loss by reconstructing the learnable high-resolution state and constraining system transition simultaneously. Specifically, RPLPO combines an encoding module for reconstructing learnable high-resolution states with a transition module for predicting future states. The two modules are jointly trained by data and PDE loss. We conduct experiments in various physical systems to demonstrate that RPLPO has significant improvement in generalization, even when observation is sparse, irregular, noisy, and PDE is inaccurate. The code is available on GitHub: RPLPO.
@@ Introduction
Using machine learning methods to model physical systems has become a promising direction in science and engineering, e.g., fluid dynamics, diffusion, and so on (Zhang et al. 2023; Wang et al. 2023). Since the data collection procedure is always both time and cost-consuming, a limited amount of data will hurt the generalization of the physical system modeling. Recently, some works have introduced PDE loss in training models and achieved promising performance (Li et al. 2021b; Gao, Sun, and Wang 2021a; Goswami et al. 2022a; Ren et al. 2023b; Huang et al. 2023) to reduce the use of costly data and improve generalization capacity of physical system models. Most works rely on high-resolution states to calculate or approximate the derivative for PDE loss. However, due to sensor limitations (Iakovlev, Heinonen, and Lähdesmäki 2020; Lütjens et al. 2022; Yin et al. 2022; Boussif et al. 2022; Iakovlev, Heinonen, and Lähdesmäki 2023), the application of PDE loss for modeling physical systems is significantly challenged by the partially observable nature of measurement. Computing partial derivatives in PDE loss with partial observation introduces significant biases, leading to a dilemma: either we avoid using PDE loss in the model, thereby weakening its generalization, or we use the biased PDE loss, which can also hurt the model’s performance. As a result, a scientific problem has been emerged:
@@
We briefly discuss the highly related works here and leave the detailed version in Technical Appendix Related Works.
Physics-informed machine learning: More and more physics-informed machine learning methods have been proposed to learn the solutions of PDEs (Karniadakis et al. 2021), including neural operator (Lu, Jin, and Karniadakis 2019; Li et al. 2020a, b, 2021b; Wang, Wang, and Perdikaris 2021; Gupta, Xiao, and Bogdan 2021; Goswami et al. 2022b; Boussif et al. 2022; Yin et al. 2023; Iakovlev, Heinonen, and Lähdesmäki 2023; Hansen et al. 2023) and physics-informed neural networks (PINNs) (Raissi, Perdikaris, and Karniadakis 2019; Yang, Meng, and Karniadakis 2021; Cai et al. 2021; Karniadakis et al. 2021). These works leveraged PDE loss to model or solve PDEs dynamics. The key to their success is the incorporation of accurate PDE, including high-resolution states or formulas of PDEs. However, they cannot be applied to learn partial observation directly, as their physics-informed training manner has significant bias when calculating by partial observation.
High-resolution state reconstruction: In computer vision (CV) and physical systems modeling, several works aim to reconstruct a high-resolution state from its partial observation, also known as super-resolution (Zhu et al. 2020; Li et al. 2021a). In CV, numerous deep learning-based models (Dong et al. 2014; Lim et al. 2017; Soh et al. 2019; Nazeri, Thasarathan, and Ebrahimi 2019; Zhao et al. 2020), and generative models (Ledig et al. 2017; Liu, Siu, and Wang 2021; Gao et al. 2023) have emerged. In physical systems, the task has attracted more and more attention (Ren et al. 2023a), which aims to use PDE loss in models (Wang et al. 2020a; Esmaeilzadeh et al. 2020; Fathi et al. 2020; Jangid et al. 2022; Ren et al. 2023b; Shu, Li, and Farimani 2023) or not need for data (Gao, Sun, and Wang 2021b; Kelshaw, Rigas, and Magri 2022; Zayats et al. 2022). Most of these works rely on high-resolution states to do supervised learning; only a small part relies on PDE loss, thus resulting in relatively larger reconstruction errors without leveraging data information or hard to model the unsteady dynamics (Gao, Sun, and Wang 2021b). There is a work (Rao et al. 2023) that is most related to our work. In this work, Chengpeng Rao, et al. also considered the high-resolution state reconstruction and the time-series prediction problem, but unlike our work, it is similar to the neural PDE solvers like PINNs rather than operator learning. It cannot generalize to different initial conditions. We have also leveraged this work as one of our baselines.
@@
Problem setting: We aim to use PDE loss to improve model’s generalization for predicting future partially observed states under partial observation in physical systems. Denote $u_{t}\triangleq u(t)\in\mathcal{U}$ as an observed state, we want to infer $u_{t+\tau}$ for $\tau>0$ at any time $t$. $\mathcal{U}$ is the functional space of form $\Omega\rightarrow\mathbb{R}^{n}$, where $\Omega\subset\mathbb{R}^{p}$ is the set of observational point location, and $n$ is the number of system variable. That is to say, $u_{t}$ is a function of $x\in\Omega$, with vectorical output $u_{t}(x)\in\mathbb{R}^{n}$; cf. examples of Section Experiments Setup. In such problems, trajectories share the same dynamics but vary by their initial conditions (ICs) $u_{0}\in\mathcal{U}$. We observe a finite training set of trajectories $\mathcal{D}$ with label and $\mathcal{B}$ without label from trajectories given different ICs $u_{0}$, using a partial spatial observation grid $\mathcal{X}\subset\Omega$ on discrete times $t\in\mathcal{T}\subset[0,T]$. In inference, the partial observation dataset is observed with $\mathcal{X}$. Note that inference is performed on test data observed from trajectories given different ICs to verify model’s generalization to different trajectories.
@@
Figure 1: RPLPO. Training (left): In the base-training period, the encoding module is trained by $\mathcal{L}_{D}$ and $\mathcal{L}_{P}^{\theta}$, and the transition module is trained by $\mathcal{L}_{D}$ and $\mathcal{L}_{P}^{\phi}$. These losses are calculated on labeled dataset $\mathcal{D}$. Then, in the two-stage fine-tuning period, the transition module is tuned by $\mathcal{L}_{P}^{\phi}$, calculated on unlabeled dataset $\mathcal{B}$, and the encoding module is tuned by $\mathcal{L}_{D}$, calculated on $\mathcal{D}$, in order. Inference (right): given the partial observation to predict future partially observed states.
@@
In our current implementation, the reason for employing U-Net is the encoding module aims to extract features from partial observation for high-resolution reconstruction, and the U-Net (Ronneberger, Fischer, and Brox 2015) is a proven model well-suited for such a task. For the transition module, we leverage FNO (Li et al. 2020b) to capture continuous and global spatial features. Note that above choices are specific to our experiments’ requirements. Future work could explore the efficacy of alternative network architectures.
@@
Ablation study on hyperparameters: We examine three critical hyperparameters in RPLPO for robustness: (1) PDE loss weight $\gamma$ from Eqn. 5, (2) lengths of recent temporal observations $n$ as input, and (3) the number of epochs $m_{1},m_{2}$ and gaps $q$ between each fine-tuning period. (i) Figure 3 (left) indicates clearer model improvement at $\gamma$ = 1E-1, chosen for this work. RPLPO shows consistent enhancement regardless of $\gamma$, demonstrating robustness. (ii) Longer $n$ values ($n=4,6$) lead to better RPLPO performance, as seen in Figure 3 (middle), by utilizing more temporal observations. (iii) Figure 3 (right) shows RPLPO’s improvements across all parameters, confirming RPLPO can adapt to varying training setups.

### v3

@@ Abstract
In science and engineering, machine learning techniques are increasingly successful in physical systems modeling (predicting future states of physical systems). Effectively integrating PDE loss as a constraint of system transition can improve the model’s prediction by overcoming generalization issues due to data scarcity, especially when data acquisition is costly. However, in many real-world scenarios, due to sensor limitations, the data we can obtain is often only partial observation, making the calculation of PDE loss seem to be infeasible, as the PDE loss heavily relies on high-resolution states. We carefully study this problem and propose a novel framework named Re-enable PDE Loss under Partial Observation (RPLPO). The key idea is that although enabling PDE loss to constrain system transition solely is infeasible, we can re-enable PDE loss by reconstructing the learnable high-resolution state and constraining system transition simultaneously. Specifically, RPLPO combines an encoding module for reconstructing learnable high-resolution states with a transition module for predicting future states. The two modules are jointly trained by data and PDE loss. We conduct experiments in various physical systems to demonstrate that RPLPO has significant improvement in generalization, even when observation is sparse, irregular, noisy, and PDE is inaccurate. The code is available on RPLPO.
@@ Introduction
Using machine learning methods to model physical systems has become a promising direction in science and engineering, e.g., fluid dynamics, diffusion, and so on (An, Kim, and James 2008; Zhang et al. 2023; Wang et al. 2023). Since the data collection procedure is always both time and cost-consuming, a limited amount of data will hurt the generalization of the physical system modeling. Recently, some works have introduced PDE loss in training models and achieved promising performance (Li et al. 2021b; Gao, Sun, and Wang 2021a; Goswami et al. 2022a; Ren et al. 2023b; Huang et al. 2023) to reduce the use of costly data and improve generalization capacity of physical system models. Most works rely on high-resolution states to calculate or approximate the derivative for PDE loss. However, due to sensor limitations (Iakovlev, Heinonen, and Lähdesmäki 2020; Lütjens et al. 2022; Yin et al. 2022; Boussif et al. 2022; Iakovlev, Heinonen, and Lähdesmäki 2023), the application of PDE loss for modeling physical systems is significantly challenged by the partially observable nature of measurement. Computing partial derivatives in PDE loss with partial observation introduces significant biases, leading to a dilemma: either we avoid using PDE loss in the model, thereby weakening its generalization, or we use the biased PDE loss, which can also hurt the model’s performance. As a result, a scientific problem has been emerged:
@@
We briefly discuss the highly related works here and leave the detailed version in Technical Appendix.
Physics-informed machine learning: More and more physics-informed machine learning methods have been proposed to learn the solutions of PDEs (Karniadakis et al. 2021), including neural operator (Lu, Jin, and Karniadakis 2019; Li et al. 2020a, b, 2021b; Wang, Wang, and Perdikaris 2021; Gupta, Xiao, and Bogdan 2021; Goswami et al. 2022b; Boussif et al. 2022; Yin et al. 2023; Iakovlev, Heinonen, and Lähdesmäki 2023; Hansen et al. 2023; Chen et al. 2023) and physics-informed neural networks (PINNs) (Raissi, Perdikaris, and Karniadakis 2019; Yang, Meng, and Karniadakis 2021; Cai et al. 2021; Karniadakis et al. 2021). These works used PDE loss to model or solve PDEs dynamics. The key to their success is the incorporation of accurate PDE, including high-resolution states or formulas. However, they cannot be applied to learn partial observation directly, as their physics-informed training manner has significant bias when calculating by partial observation.
High-resolution state reconstruction: In computer vision (CV) and physical systems modeling, several works aim to reconstruct a high-resolution state from its partial observation, also known as super-resolution (Zhu et al. 2020; Li et al. 2021a). In CV, some deep learning models (Dong et al. 2014; Lim et al. 2017; Soh et al. 2019; Nazeri, Thasarathan, and Ebrahimi 2019; Zhao et al. 2020), and generative models (Ledig et al. 2017; Liu, Siu, and Wang 2021; Gao et al. 2023) have emerged. In physical systems, the task has attracted more and more attention (Ren et al. 2023a), which aims to use PDE loss in models (Wang et al. 2020a; Esmaeilzadeh et al. 2020; Fathi et al. 2020; Jangid et al. 2022; Ren et al. 2023b; Shu, Li, and Farimani 2023) or not need for data (Gao, Sun, and Wang 2021b; Kelshaw, Rigas, and Magri 2022; Zayats et al. 2022). Most of these works rely on high-resolution states to do supervised learning; only a small part relies on PDE loss, thus resulting in relatively larger reconstruction errors without leveraging data information or hard to model the unsteady dynamics (Gao, Sun, and Wang 2021b). There is a work (Rao et al. 2023) that is most related to our work. In this work, Chengpeng Rao, et al. also considered the high-resolution state reconstruction and the time-series prediction problem, but unlike our work, it is similar to the neural PDE solvers like PINNs rather than operator learning. It cannot generalize to different initial conditions. We have also leveraged this work as one of our baselines.
@@
Problem setting: We aim to use PDE loss to improve model’s generalization for predicting future partially observed states under partial observation in physical systems. Denote $u_{t}\triangleq u(t)\in\mathcal{U}$ as an observed state, we want to infer $u_{t+\tau}$ for $\tau>0$ at any time $t$. $\mathcal{U}$ is the functional space of form $\Omega\rightarrow\mathbb{R}^{n}$, where $\Omega\subset\mathbb{R}^{p}$ is the set of observational point location, and $n$ is the number of system variable. That is to say, $u_{t}$ is a function of $x\in\Omega$, with vectorical output $u_{t}(x)\in\mathbb{R}^{n}$; cf. examples of Section Experiments Setup. In such problems, trajectories share the same dynamics but vary by their initial conditions (ICs) $u_{0}\in\mathcal{U}$. We observe a finite training set of trajectories $\mathcal{D}$ with label and $\mathcal{B}$ lacks future observations (without label), using a partial spatial observation grid $\mathcal{X}\subset\Omega$ on discrete times $t\in\mathcal{T}\subset[0,T]$. In inference, the partial observation dataset is observed with $\mathcal{X}$. Note that inference is performed on test data observed from trajectories given different ICs to verify model’s generalization to different trajectories.
@@
Figure 1: RPLPO. Training (left): In the base-training period, the encoding module is trained by $\mathcal{L}_{D}$ and $\mathcal{L}_{P}^{\theta}$, and the transition module is trained by $\mathcal{L}_{D}$ and $\mathcal{L}_{P}^{\phi}$. These losses are calculated on labeled dataset $\mathcal{D}$. Then, in the two-stage fine-tuning period, the transition module is tuned by $\mathcal{L}_{P}^{\phi}$, calculated on unlabeled dataset $\mathcal{B}$, and the encoding module is tuned by $\mathcal{L}_{D}$, calculated on $\mathcal{D}$, in order. Inference (right): given the partial observation to predict future partially observed states. For the two-stage fine-tuning period, please check Technical Appendix Figure 4.
@@
Ablation study on hyperparameters: We examine three critical hyperparameters in RPLPO for robustness: (1) PDE loss weight $\gamma$ from Eqn. 5, (2) lengths of recent temporal observations $n$ as input, and (3) the number of epochs $m_{1},m_{2}$ and gaps $q$ between each fine-tuning period. (i) Figure 3 (left) indicates clearer model improvement at $\gamma$ = 1E-1, chosen for this work. RPLPO shows consistent enhancement regardless of $\gamma$, demonstrating robustness. (ii) Longer $n$ values ($n=4,6$) lead to better RPLPO performance, as seen in Figure 3 (middle), by utilizing more temporal observations. (iii) Figure 3 (right) shows RPLPO’s improvements across all parameters, confirming RPLPO can adapt to varying training setups. We also study the impact of two PDE losses with different $\gamma$ and the higher upscale factor of the encoding module. Their results are shown in Technical Appendix Table 18 and Table 24.
@@ Conclusion
Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC NO. 12302365).